OpenStudy (anonymous):

given the differential equation \[\left( 1-x^{2} \right) dx/dy=2y \] , do you have to use partial fraction expansion to solve, and what is the general solution?

OpenStudy (jamesj):

This equal is very clearly separable and when you do that, I don't see any terms in the denominator, do you?

OpenStudy (jamesj):

*equation

OpenStudy (anonymous):

i separated it, and i got \[dy/y=2/(1-x ^{2})dx\] and from there im stuck

OpenStudy (anonymous):

i dont know how to integrate the right side

OpenStudy (wasiqss):

its easy , u can split into partial fraction (1+x)(1-x)

OpenStudy (wasiqss):

then its easy

OpenStudy (jamesj):

As you've written the equation originally, you have \[ \int (1-x^2) \ dx = \int 2y \ dy \]

OpenStudy (jamesj):

Did you invert the derivative in your original expression?

OpenStudy (wasiqss):

yehh james rite ,

OpenStudy (anonymous):

oops. yes i did

OpenStudy (anonymous):

im trying to figure out how to split 2/(1-x^2) into partial fractions but i forgot the process

OpenStudy (wasiqss):

man u made wrong equations, james made it rite for u

OpenStudy (anonymous):

i figured it out. partial fraction expansion gives me 1/1-x + 1/1-x